# Co-Designing Multimodal Systems for Accessible Asynchronous Dance Instruction

### v3

@@ 5.2.2. Application Across Dance Styles
Dance encompasses a wide range of styles differing in movement standardization and learning techniques (Conroy and Bresnahan, 2025; Foster, 1986; Nicholas and Morris, 2004). Certain styles emphasize standard positions and sequences (e.g., ballet (Bennett, 2008), Indian classical dance (Puri, 2004)). Other styles like Butoh (Kasai, 1999), Gaga (Katan-Schmid, 2016), and Contemporary Improvisation (Nakano and Okada, 2012) prioritize sensation, and internal imagery over the reproduction of fixed shapes (Lepecki, 2006). Instruction for these dance styles rely on sensory descriptions to guide dancers (Seham and and, 2015), offering more inclusive learning by de-emphasizing the reliance on visual replication. In this work, we address asynchrnous learning from online dance tutorial videos, which predominantly feature structured movement routines and timing. This sequence-based, visually taught format presents distinct challenges for BLV audiences. However, we acknowledge the broader landscape of more accessible, improvisational dance practices that are less represented in such online dance tutorials.
By examining contrasts between groups, our workshops revealed differences in modality preferences and system designs across different dance styles. Tactile and non-verbal communication are critical aspects of partner dancing (Grammer et al., 1998), aligning with our findings where haptics was favored by all groups addressing partner dance clips (DI4, DI7, DI10). Haptic feedback is less descriptive than sounds and narration, and was therefore less suited for conveying complex movements in expressive dance instruction systems (e.g., for Contemporary dance). In such cases, sound cues and tactile metaphors (DI5) were more beneficial. For dance styles with exact, pre-defined movement where precision is important (e.g., for Ballet), DI2-DI3 are more relevant. More broadly, future systems should be customized for individual dance styles (DI14) rather than assume a single system will suit all dances.
@@ 5.3. Limitations
Our workshop focused on multi-modal technologies for asynchronous dance education, and the specific dance styles considered were dependent on the background of the dance teachers within each group. Although our study found significant differences in system preferences across different dance styles, we did not consider all dance styles and future work with a narrowed focus on a specific style would provide more narrowed considerations. Further, our insights may fail to support asynchronous learning of dance practices that are highly improvisational in which learning movements or sequences of movements from others may not be prioritized. Additionally, our workshops are exploratory and thus our design implications, system designs, and modality considerations are preliminary. Future work may explore empirical validation of design implications and initial low-fidelity systems through high-fidelity prototype development, testing, and refinement with co-design. During our workshops, we provided participants with low-fidelity prototyping materials (e.g., a simple soundboard, tactile stickers to represent haptics), but more advanced haptic devices, spatial audio systems, and automated movement analysis may expand the scope of design possibilities and limitations considered by the participants. Furthermore, the in-person setting does not fully reflect the experience of BLV individuals learning asynchronously. Future evaluation of the design implications and proposed system workflows in remote settings would be beneficial to validate or refine our design implications.
@@ 5.4. Future Work
Our work reveals future opportunities for expanding asynchronous instruction to new types of movements and new learners, exploring involving instructors in asynchronous instruction systems, and enabling long-term adaptive asynchronous instruction systems.
@@
Multimodal instruction can generalize across movement practices, abilities, and learning preferences. Although our work focuses on dance, the design implications we propose extend to other domains where movement is taught through routines composed of specific poses and transitions. For example, martial arts (Luft et al., 2023; Jennings et al., 2020) and aerobics (Sun, 2021) rely on choreographed sequences that combine discrete positions with fluid movements. Similar to dance, these practices require learners to internalize not just static poses but timing, rhythm, and expressiveness of transitions. Our findings suggest that multimodal vocabularies, layered feedback, and staged practice could support instruction in these contexts as well. For instance, using existing or system-specific named poses could form a multimodal vocabulary for martial arts routines, while haptic cues could guide stance corrections or foot placement.
Non-verbal sound mappings could convey qualities such as speed, impact, or balance, paralleling how they are used to express weight and fluidity in dance. Extending to other populations also reveals new opportunities for future work. For DeafBlind learners, systems could limit reliance on auditory feedback while still supporting rhythmic and expressive qualities through haptics. Furthermore, the decomposition of routines and movements to reduce cognitive load may be beneficial for neurodivergent populations (Stavraki, [n. d.]; Sharer et al., 2016).
@@
While future systems may be fully automated (e.g., to convert a dance video to an instructional system, similar to Vid2Coach for cooking videos (Huh et al., 2025)), dance instructors revealed that they may be interested in making custom material for systems (e.g., D6 mentioned they would want to re-make a dance video with useful descriptions for BLV audiences). In the future, we may explore authoring systems for instructors seeking to create accessible asynchronous instruction materials. For example, the instructors could define their own movement descriptions, movement names, tactile cues, and haptic cues that could be flexibly applied by the system as students learn. In addition, with instructor involement, future work can explore opportunities to blend synchronous and asynchronous approaches (e.g., automatically generating an instructional practice system based on an in-class activity).
@@
While our findings clearly reveal that BLV learners want to be able to customize their system experiences, this requires BLV learners knowing what options they want to use up front. In the future, we could explore closed-loop approaches to automatically adjust the descriptiveness level, modality level, or feedback type (e.g., positive vs. corrective) based on the current performance of the learner. In the best case, learners may be able to adjust their customization preferences, then let the system adapt within their preferences to achieve a better outcome compared to what learner-driven customization or system-driven adaptivity could do alone. To co-design such highly technical systems, future work may explore Wizard of Oz instructional systems over a longer instruction period to surface learner preferences.
@@ 6. Conclusion
Accessible body movement education as a whole is an underexplored area of research in HCI. We conducted three co-design workshops with BLV individuals, dance instructors, and domain-specific experts (i.e., sound, haptics, AD) to understand user needs and preferences for asynchronous dance instruction systems. We identified the challenges BLV learners face when learning dance with videos, and we examined how alternative modalities—verbal, sound, haptic, and tactile—can be orchestrated across stages of learning to provide feedback-oriented, customized instruction. From these findings, we surfaced co-designed system architectures and developed a set of design implications (DI1–DI14) that address instruction of movement structure, timing, and expressiveness and provide feedback, practice, and customization. From the insights gained in this work, we can inform the design of systems addressing body movement education for a diverse set of populations and activities involving sequences of movements, movement expression, and precise timing.

### v4

@@ 5.2.2. Application Across Dance Styles
Dance encompasses a wide range of styles differing in movement standardization and learning techniques (Conroy and Bresnahan, 2025; Foster, 1986; Nicholas and Morris, 2013). Certain styles emphasize standard positions and sequences (e.g., ballet (Bennett, 2008), Indian classical dance (Puri, 2004)). Other styles like Butoh (Kasai, 1999), Gaga (Katan-Schmid, 2016), and Contemporary Improvisation (Nakano and Okada, 2012) prioritize sensation, and internal imagery over the reproduction of fixed shapes (Lepecki, 2006). Instruction for these dance styles rely on sensory descriptions to guide dancers (Seham and and, 2015), offering more inclusive learning by de-emphasizing the reliance on visual replication. In this work, we address asynchronous learning from online dance tutorial videos, which predominantly feature structured movement routines and timing. This sequence-based, visually taught format presents distinct challenges for BLV audiences. However, we acknowledge the broader landscape of more accessible, improvisational dance practices that are less represented in such online dance tutorials.
By examining contrasts between groups, our workshops revealed differences in modality preferences and system designs across different dance styles (DI14). Tactile and non-verbal communication are critical aspects of partner dancing (Kinoe and Sato, 2022), aligning with our findings where haptics was favored by all groups addressing partner dance clips (DI4, DI7, DI10). Haptic feedback is less descriptive than sounds and narration, and was therefore less suited for conveying complex movements in expressive dance instruction systems (e.g., for Contemporary dance). In such cases, sound cues and tactile metaphors (DI5) were more beneficial. For dance styles with exact, pre-defined movement where precision is important (e.g., for Ballet), detailed verbal descriptions (DI2-DI3) were more critical.
@@ 5.3. Limitations
Our workshop focused on multimodal technologies for asynchronous dance instruction, and the specific dance styles considered were dependent on the background of the dance teachers within each group. Although our study found significant differences in system preferences across different dance styles, we did not consider all dance styles and future work with a narrowed focus on a specific style would provide more narrowed considerations. Further, our insights may fail to support asynchronous learning of dance practices that are highly improvisational in which learning movements or sequences of movements from others may not be prioritized. Additionally, our workshops are exploratory and thus our design implications, system designs, and modality considerations are preliminary. Future work may explore empirical validation of design implications and initial low-fidelity systems through high-fidelity prototype development, testing, and refinement with co-design. During our workshops, we provided participants with low-fidelity prototyping materials (e.g., a simple soundboard, tactile stickers to represent haptics), but more advanced haptic devices, spatial audio systems, and automated movement analysis may expand the scope of design possibilities and limitations considered by the participants. Furthermore, the in-person setting does not fully reflect the experience of BLV individuals learning asynchronously. Future evaluation of the design implications and proposed system workflows in remote settings would be beneficial to validate or refine our design implications.
@@ 5.4. Future Work
Our work reveals future opportunities for expanding asynchronous instruction to new types of movements and learners, involving instructors in asynchronous instruction systems, and enabling long-term adaptive asynchronous instruction systems.
@@
Multimodal instruction can generalize across movement practices, abilities, and learning preferences. Although our work focuses on dance, the design implications we propose extend to other domains where movement is taught through routines composed of specific poses and transitions. For example, martial arts (Luft et al., 2023; Jennings et al., 2020) and aerobics (Sun, 2021) rely on choreographed sequences that combine discrete positions with fluid movements. Similar to dance, these practices require learners to internalize not just static poses but timing, rhythm, and fluidity. Our DIs of multimodal vocabularies and feedback could support instruction in these contexts as well. For instance, using existing or system-specific named poses could form a multimodal vocabulary for martial arts routines, while haptic cues guide stance corrections or foot placement.
Sound mappings could convey qualities such as speed, impact, or balance, paralleling how they are used to express weight and fluidity in dance. Extending to other populations also reveals new opportunities for future work. For DeafBlind learners, systems could limit reliance on auditory feedback and emphasize using haptics to convey rhythm and expressiveness. Furthermore, the segmentation of routines and movements into clear steps may be beneficial for neurodivergent populations (Stavraki, [n. d.]; Sharer et al., 2016).
@@
While future systems may be fully automated (e.g., converting a dance video to an instructional system, similar to Vid2Coach for cooking videos (Huh et al., 2025)), dance instructors revealed that they may be interested in making custom material for systems (e.g., D6 expressed wanting to re-make dance videos with useful descriptions for BLV audiences). In the future, we may explore authoring systems for instructors seeking to create accessible asynchronous instruction materials. For example, the instructors could define their own movement descriptions, names, sounds, and haptic cues that are incrementally applied by the system as students learn. In addition, with instructor involvement, future work can explore opportunities to blend synchronous and asynchronous approaches (e.g., generating an instructional practice system based on in-class activities).
@@
While our findings clearly reveal that BLV learners want to be able to customize their system experiences, this requires BLV learners knowing what options they want to use up front. In the future, we could explore closed-loop approaches to automatically adjust the descriptiveness level, modalities, or feedback type (e.g., positive vs. corrective) based on learner performance. Learners may be able to adjust their customization preferences, then let the system adapt within their preferences to achieve a better outcome compared to what learner-driven customization or system-driven adaptivity could do alone. To co-design such highly technical systems, future work may explore Wizard of Oz instructional systems over a longer instruction period to surface learner preferences.
@@ 6. Conclusion
Accessible movement education as a whole is underexplored in HCI. We conducted three co-design workshops with BLV individuals, dance instructors, and domain-specific experts (sound, haptics, AD) to understand user needs and preferences for asynchronous dance instruction systems. We identified the challenges BLV learners face when learning dance with videos, and examined how different modalities can be applied for learning and feedback. From these findings, we surfaced co-designed system architectures and developed a set of design implications (DI1–DI14) that address instruction of movement structure, timing, and expressiveness and provide feedback, practice, and customization. From the insights gained in this work, we can inform the design of systems addressing movement education for a diverse set of populations and activities involving movement sequences, expressiveness, and precise timing.
Acknowledgements.
We thank our participants for their contribution in our workshops.